OpenStudy (anonymous):

What's the integral of: sqrt(8x-x^2)?

OpenStudy (anonymous):

i think this is a set up for completing the square

OpenStudy (anonymous):

I've already did that, but I cannot come to the final answer at all.

OpenStudy (anonymous):

done*

OpenStudy (anonymous):

you got to \[\int \sqrt{16-(x-4)^2}dx\]right?

OpenStudy (anonymous):

Yep! But I can't get to the final part. Used trig substitution and everything.

OpenStudy (anonymous):

one thing you could do is cheat and look in the back of your book, where no doubt it has a formula for \[\int \sqrt{a^2-u^2}du\]

OpenStudy (anonymous):

the other way i guess it to make the substitution \(u=4\sin(\theta)\)

OpenStudy (anonymous):

that is after you already make the mental substitution \(u=x-1\) with \(du=dx\) so there is nothing really to change

OpenStudy (anonymous):

This is what I got for the answer. \[8\arcsin((x-4)/4) + (x-4)\sqrt{8x-x^2}\] Also, this is online, so no book.

OpenStudy (anonymous):

ok lets see what we get

OpenStudy (anonymous):

\[u=4\sin(\theta), du = 4\cos(\theta)d\theta\] and so after the magic disapperance of the radical you get \[\int 16\cos^2(\theta)d\theta\]

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

out pops the 16 and you rewrite this mess at \[8\int (\cos(2\theta)+1)d\theta\]

OpenStudy (anonymous):

Hold up. Why not just do integration by parts on the cos^2?

OpenStudy (anonymous):

how is that going to work? well, it might work, but you are going to get a mess i think standard way to integrate cosine squared is like that

OpenStudy (anonymous):

Alright. We weren't taught it like that, so please, continue.

OpenStudy (anonymous):

we can try that if you like you are going to do one of those integrals around in a circle i think way too much work

OpenStudy (anonymous):

in any case if we do this the answer becomes \[4\sin(2\theta)+8\theta\]

OpenStudy (anonymous):

then put \(\theta=\sin^{-1}(\frac{x-4}{4})\) to get the answer i think

OpenStudy (anonymous):

Ok. I'm gonna try that real quick.

OpenStudy (anonymous):

actually your answer looks good to me in any case

OpenStudy (anonymous):

same answer i think that i get , matter of fact i am almost sure of it

OpenStudy (anonymous):

Odd. It's saying that it's not the correct answer, but that there's always more than one possibility.

OpenStudy (anonymous):

yeah i hate that nonsense you can probably write this in a raft of different ways, all of which are the same, or that vary by a constant

OpenStudy (anonymous):

Alright. Well, honestly, I like your way of doing it much, much better. Far less clutter and BS to deal with.

OpenStudy (anonymous):

i am going to make a bet my best is that if you put \(\int \sqrt{8x-x^2}dx\) in to wolfram, and then put in \(\int \sqrt{16-(x-4)^2}dx\) you get two different answers, even though they are identical

OpenStudy (anonymous):

I'll try that now.

OpenStudy (anonymous):

Online math is the worst.

OpenStudy (anonymous):

Alright. Just got it. I rewrote a previous way I did it, just to see, and it accepted it. \[8\arcsin((x-4)/4) + ((2x - 8) / 4)\sqrt(8x-x^2) + c\]

OpenStudy (anonymous):

don't you love on line courses?

OpenStudy (anonymous):

Oh, they truly are a blessing to college students everywhere.

OpenStudy (anonymous):

mylabsplus webassign thinkwell?

OpenStudy (anonymous):

WebWork, which I think is the same as WebAssign.

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

yw